# ConSiDERS-The-Human Evaluation Framework: Rethinking Human Evaluation for Generative Large Language Models

### v1

@@ 1 Introduction
In this position paper, first, we argue that to design and interpret the results of human evaluation accurately, the evaluation pipeline needs to be human-centric in the age of generative AI, accounting for human evaluators and their cognitive biases. The field of user experience (UX) takes into account the emotional states of a user, a.k.a. how a user feels Marques et al. (2021); Hartson and Pyla (2012). It is a well-known fact in UX that users tend to be heavily influenced by aesthetic aspects, while actual function or usability aspects take a second place when users perceive a system as useful, leading to the notion “what is beautiful is usefulSonderegger and Sauer (2010); Tuch et al. (2012); Hamborg et al. (2014). Aesthetics also extends to language, factors such as fluency can affect the evaluation, outweighing the actual content or substance Reber (2011). Studies in human-computer interface (HCI), cognitive, and social psychology have demonstrated that processing fluency – the ease with which information is perceived and processed in the human mind – has a positive effect on evaluation Preßler et al. (2023); Tsai and Thomas (2011); Greifeneder and Bless (2017). Current state-of-the-art (SOTA) LLMs tend to be quite fluent & produce content that is easy to read & understand, as a result, users can conflate fluency and usefulness. Therefore, we need to closely examine our human evaluation before reaching conclusions such as – The LLM can perform function <x> similar to or better than a trained professional. NLP evaluation procedures, therefore, at the very least must delineate style vs. substance.
Secondly, the effectiveness of the test set in measuring the capabilities of a model is critical, as ineffective test sets cannot adequately evaluate these models, a common theme that has surfaced in many leader-boards and public data sets Tedeschi et al. (2023); Elangovan et al. (2021)
@@
1) We propose a framework, a structure for organizing and contextualizing human evaluation, that can be customized and adapted to specific contexts. Our proposed framework – the ConSiDERS-The-Human evaluation framework, has 6 pillars:
@@ 2.3 Cognitive biases
Cognitive biases are systematic errors in human judgment or aspects that drive irrational behavior Ellis (2018). This is usually a result of using heuristics to quickly make a decision, and there are several types of cognitive biases that can be broadly categorized into a) too much information b) lack of information or lack of understanding or meaning associated with the information c) need to act or make judgements fast d) information that is remembered or recalled Azzopardi (2021). As an example, when a user is presented with a long list (too much information) during information search and retrieval, to quickly filter out information the user would simply click on the first link demonstrating position bias Azzopardi (2021).
@@ 3.1 User Experience
Processing fluency – the ease with which information is processed by the human mind Tsai and Thomas (2011) – affect factors such as perceived truthfulness and usefulness of statements. These lessons from psychology also apply to NLG human evaluation, where the human evaluation strategy needs to isolate the effects of linguistic fluency vs. aspects such as factual correctness. As presented in the introduction section, the notion “what is beautiful is usefulSonderegger and Sauer (2010); Hamborg et al. (2014) also extends to language, where information that is presented in an easy-to-process manner can be perceived as true Schwarz (2006). To be clear, we are not calling for linguistic fluency & coherence to be trivialized, on the contrary, we highlight how powerful its influence is on human judgment and evaluation. In the following section 3.1.1, we provide a few tips on how to isolate such effects in NLG evaluation.
@@ 3.1.1 Recommendations to mitigate common cognitive biases in NLG evaluation
Mitigation: Tasks such as fact-checking and factual completeness that require inspecting individual traits must be ideally split into atomic facts as detailed in Section 3.2 to isolate the impact of fluency & ease of information processing vs. facts.
@@ 3.4 Differentiating
To differentiate is to identify the differences between thingsvocabulary.com . The test sets used in evaluation must be able to differentiate between the various capabilities as well as the weaknesses of generative LLMs. For instance, when the test sets are relatively easy, most models can achieve high scores and seem very capable. Conversely, when the test sets are very difficult models might achieve low scores, making the models seem ineffective. Therefore, tests that are ineffective in differentiating capabilities can lead to (a) incorrect conclusions about the capabilities of a model (b) poor calibration or ranking of various models (c) inability to identify any improvements or degradation between model versions, despite spending significant resources retraining new models. Hence, constructing effective tests that can differentiate model capabilities is crucial, otherwise evaluation simply results in wasted effort.
@@
Traditionally, models have been evaluated using public datasets and benchmark such as GLUE (Wang et al., 2019b) and SuperGLUE (Wang et al., 2019a). These evaluations are not without their share of problems as models can exploit weaknesses in the official test sets relying on shortcuts or spurious correlation – such as length of the input to predict the target label achieving high performance yet non-generalizable beyond the official test set McCoy et al. (2019); Elangovan et al. (2023); Gururangan et al. (2018). In addition to these problems, evaluating the current generation of SOTA LLMs poses further challenges. 1) LLMs are trained on billions of tokens available on the internet, and the training data used is rarely well documented. Hence, these LLMs may have consumed public benchmark test sets as part of their training data Magar and Schwartz (2022); Sainz et al. (2023). 2) The models are generative, producing natural language output unlike a model trained on a classification task, making it difficult to automate evaluation and thus are far more reliant on human evaluation. 3) The LLMs are capable of following natural language instructions to solve a vast variety of tasks, hence they need to be evaluated on a range of instructions provided as prompts that emulate end-user use cases.
@@
Emulating end-user use cases is crucial, as there are fundamental differences in the tasks in NLP benchmarks used to evaluate a model vs. the kind of questions or problems the users ask an LLM like ChatGPT. Firstly, end users tend to be very creative and prompt the LLM for all kinds of queries. Secondly, the same message or question can be framed (prompted) in many ways. Hence, test scenarios need to consider the various natural ways in which users write their intentions when interacting with a LLM. If these aspects are not taken account when constructing a test set, solely relying on traditional NLP benchmark tasks such as sentiment analysis or linguistic capability can result in traditional SOTA models outperforming ChatGPT Kocoń et al. (2023), when clearly this is not reflective of the end-user experience. Hence, we argue that the test cases used for evaluating LLMs need to represent end user scenarios, in addition to standard NLP tasks. Benchmarks such as Big-Bench bench authors (2023) take a step towards this direction by enabling GitHub contributors to add new tasks to the benchmark. Furthermore, the same prompts may not be effective against all LLMs, hence these prompts might often have to be customized for individual models, making benchmarking non-trivial.
@@ 3.5 Responsible
In human evaluation, the Responsible pillar has to consider two aspects, a) Is the model behavior responsible? b) Do the human evaluators introduce bias to the evaluation results? While there is no formal definition of Responsible AI, at the very least it entails fairness, safety, truthfulness, and privacy Sun et al. (2024). Papers with “responsible” in their title results in 20 papers, expanding the search to include terms such as “bias” and “privacy” results in $\approx$1000 or 1% of the papers in ACL Anthology, while none of the human evaluation papers mention responsible AI related keywords in their title or abstract (query in Appendix A.2.8), a telltale sign that more work is needed.
@@
Generative LLMs are susceptible to leaking private details such as person identifiable information (PII) from the training data, including when subject to privacy attacks Vakili and Dalianis (2023); Li et al. (2023). One method to mitigate such privacy leaks is to obfuscate PII information such as names and locations from training data using techniques such as Pseudonymization – recognizing privacy-sensitive information and replacing them with realistic substitute Vakili and Dalianis (2023) and differential-privacy-based approaches that add noise to the input Chen et al. (2023). Safety ensures that the models do not generate content that harms a person’s physical safety or mental health Mei et al. (2023); Rusert et al. (2022). Red teams, a group of people authorized to imitate a potential adversary’s attack or exploitation capabilities, are often used to evaluate the robustness, safety, and privacy of LLMs Perez et al. (2022); Radharapu et al. (2023).
@@ 3.6 Scalability
Human evaluation is expensive, yet large volumes of test cases are necessary to differentiate model capabilities and ensure consistency. Hence, optimizations to reduce cost and time is crucial for wider adoption. Automating parts of human evaluation can reduce cost. For instance, automation might be potentially helpful to shortlist a set of candidates for human evaluation. While the shortcomings of n-gram-based automated evaluations using metrics like Rouge Lin (2004) are well studied Deutsch et al. (2022), approaches such as using LLMs to evaluate LLMs Lin and Chen (2023); Chiang and Lee (2023) needs further exploration. Firstly, the effectiveness of LLM-based evaluation is measured using its correlation with human evaluation, hence the human evaluation procedures need to be strengthened first to draw comprehensive and robust conclusions creating a chicken-and-egg problem. Secondly, LLMs to evaluate LLMs should take into account the differences between perception-based metrics and evaluations that rely on facts, as discussed in Section 3.1.1.

### v2

@@ 1 Introduction
In this position paper, first, we argue that to design and interpret the results of human evaluation accurately, the evaluation pipeline needs to be human-centric in the age of generative AI, accounting for human evaluators and their cognitive biases. The field of user experience (UX) takes into account the emotional states of a user, a.k.a. how a user feels Marques et al. (2021); Hartson and Pyla (2012). It is a well-known fact in UX that users tend to be heavily influenced by aesthetic aspects, while actual function or usability aspects take a second place when users perceive a system as useful, leading to the notion “what is beautiful is usefulSonderegger and Sauer (2010); Tuch et al. (2012); Hamborg et al. (2014). Aesthetics also extends to language. Factors such as fluency can affect the evaluation, outweighing the actual content or substance Reber (2011). Studies in human-computer interface (HCI), cognitive, and social psychology have demonstrated that processing fluency – the ease with which information is perceived and processed in the human mind – has a positive effect on evaluation Preßler et al. (2023); Tsai and Thomas (2011); Greifeneder and Bless (2017). Current state-of-the-art (SOTA) LLMs tend to be quite fluent and produce content that is easy to read and understand, and as a result, users can conflate fluency and usefulness. Therefore, we need to closely examine our human evaluation before reaching conclusions such as – The LLM can perform function <x> similar to or better than a trained professional. NLP evaluation procedures, therefore, at the very least must delineate style vs. substance.
Secondly, the effectiveness of the test set in measuring the capabilities of a model is critical, as ineffective test sets cannot adequately evaluate these models, a common theme that has surfaced in many leader-boards and public data sets Tedeschi et al. (2023); Elangovan et al. (2021).
@@
1) We propose a framework, a structure for organizing and contextualizing human evaluation, that can be customized and adapted to specific contexts. Our proposed framework – the ConSiDERS-The-Human evaluation framework – has 6 pillars:
@@ 2.3 Cognitive biases
Cognitive biases are systematic errors in human judgment or aspects that drive irrational behavior Tversky and Kahneman (1974); Ellis (2018). This is usually a result of relying on heuristics to make a decision. There are many types of cognitive biases, the sources can be broadly categorized into a) too much information b) lack of information or lack of understanding or meaning associated with the information c) need to act or make judgments fast d) information that is remembered or recalled Azzopardi (2021). For example, when a user is presented with a long list (too much information) during information retrieval, to quickly filter out information, the user would simply click on the first link due to position bias Azzopardi (2021).
@@ 3.1 User Experience
Processing fluency – the ease with which information is processed by the human mind Tsai and Thomas (2011) – affects factors such as perceived truthfulness and usefulness of statements. These lessons from psychology also apply to NLG human evaluation, where the human evaluation strategy needs to isolate the effects of linguistic fluency vs. aspects such as factual correctness. As presented in the introduction section, the notion “what is beautiful is usefulSonderegger and Sauer (2010); Hamborg et al. (2014) also extends to language, where information that is presented in an easy-to-process manner can be perceived as true Schwarz (2006). To be clear, we are not calling for linguistic fluency and coherence to be trivialized. On the contrary, we highlight how powerful its influence is on human judgment and evaluation. In the following section 3.1.1, we provide a few tips on how to isolate such effects in NLG evaluation.
@@ 3.1.1 Recommendations to mitigate common cognitive biases in NLG evaluation
Mitigation: Tasks such as fact-checking and factual completeness that require inspecting individual traits must be ideally split into atomic facts as detailed in Section 3.2 to isolate the impact of fluency and ease of information processing vs. facts.
@@ 3.4 Differentiating
To differentiate is to identify the differences between thingsvocabulary.com . The test sets used in evaluation must be able to differentiate between the various capabilities as well as the weaknesses of generative LLMs. For instance, when the test sets are relatively easy, most models can achieve high scores and seem very capable. Conversely, when the test sets are very difficult, models might achieve low scores, making the models seem ineffective. Therefore, tests that are ineffective in differentiating capabilities can lead to (a) incorrect conclusions about the capabilities of a model, (b) poor calibration or ranking of various models, or (c) inability to identify any improvements or degradation between model versions, despite spending significant resources retraining new models. Hence, constructing effective tests that can differentiate model capabilities is crucial, otherwise evaluation simply results in wasted effort.
@@
Traditionally, models have been evaluated using public datasets and benchmark such as GLUE (Wang et al., 2019b) and SuperGLUE (Wang et al., 2019a). These evaluations are not without their share of problems as models can exploit weaknesses in the official test sets relying on shortcuts or spurious correlation – such as length of the input – to predict the target label achieving high performance yet non-generalizable beyond the official test set McCoy et al. (2019); Elangovan et al. (2023); Gururangan et al. (2018). In addition to these problems, evaluating the current generation of SOTA LLMs poses further challenges. 1) LLMs are trained on billions of tokens available on the internet, and the training data used is rarely well documented. Hence, these LLMs may have consumed public benchmark test sets as part of their training data Magar and Schwartz (2022); Sainz et al. (2023). 2) The models are generative, producing natural language output unlike a model trained on a classification task, making it difficult to automate evaluation and thus are far more reliant on human evaluation. 3) The LLMs are capable of following natural language instructions to solve a vast variety of tasks, hence they need to be evaluated on a range of instructions provided as prompts that emulate end-user use cases.
@@
Emulating end-user use cases is crucial, as there are fundamental differences in the tasks in NLP benchmarks vs. the kind of questions or problems users can ask an LLM. Firstly, end users tend to be very creative and prompt the LLM for all kinds of queries. This diversity is particularly important in safety critical applications such as Medicine, as evaluation blindspots can potentially lead to harmful consequences. Secondly, the same message or question can be framed (prompted) in many ways. Hence, test scenarios need to consider the various natural ways in which users write their intentions when interacting with a LLM. If these aspects are not taken account when constructing a test set, solely relying on traditional NLP tasks such as sentiment analysis can result in traditional SOTA models outperforming ChatGPT Kocoń et al. (2023), when clearly this is not reflective of the end-user experience. Hence, we argue that the test cases used for evaluating LLMs need to represent end user scenarios, in addition to standard NLP tasks. Benchmarks such as Big-Bench bench authors (2023) take a step towards this direction by enabling GitHub contributors to add new tasks to the benchmark. Furthermore, the same prompts may not be effective against all LLMs. Hence, prompts might often have to be customized for individual models, making benchmarking non-trivial.
@@ 3.5 Responsible
In human evaluation, the Responsible pillar has to consider two aspects, a) Is the model behavior responsible? b) Do the human evaluators introduce bias to the evaluation results? While there is no formal definition of Responsible AI, at the very least it entails fairness, safety, truthfulness, and privacy Sun et al. (2024). Searching papers with “responsible” in their title results in 20 papers, expanding the search to include terms such as “bias” and “privacy” results in $\approx$1000 or 1% of the papers in ACL Anthology, while none of the human evaluation papers mention responsible AI related keywords in their title or abstract (query in Appendix A.2.8), a telltale sign that more work is needed.
@@
Generative LLMs are susceptible to leaking private details such as person identifiable information (PII) from the training data, including when subject to privacy attacks Vakili and Dalianis (2023); Li et al. (2023). One method to mitigate such privacy leaks is to obfuscate PII information such as names and locations from training data using techniques such as Pseudonymization – recognizing privacy-sensitive information and replacing them with realistic substitute Vakili and Dalianis (2023) and differential-privacy-based approaches that add noise to the input Chen et al. (2023). Safety ensures that the models do not generate content that harms a person’s physical safety or mental health Mei et al. (2023); Rusert et al. (2022). Red teams, a group of people authorized to imitate an adversary’s attack or exploitation capabilities, are used to evaluate the robustness, safety, and privacy of LLMs Perez et al. (2022); Radharapu et al. (2023).
@@ 3.6 Scalability
Human evaluation is expensive, yet large volumes of test cases are necessary to differentiate model capabilities and ensure consistency. Hence, optimizations to reduce cost and time is crucial for wider adoption. Automating parts of human evaluation can reduce cost. For instance, automation might be potentially helpful to shortlist a set of candidates for human evaluation. While the shortcomings of n-gram-based automated evaluations using metrics like Rouge Lin (2004) are well studied Deutsch et al. (2022), approaches such as using LLMs to evaluate LLMs Lin and Chen (2023); Chiang and Lee (2023) need further exploration. Firstly, the effectiveness of LLM-based evaluation is measured using its correlation with human evaluation, hence the human evaluation procedures need to be strengthened first to draw comprehensive and robust conclusions, creating a chicken-and-egg problem. Secondly, LLMs to evaluate LLMs should take into account the differences between perception-based metrics and evaluations that rely on facts, as discussed in Section 3.1.1.